OpenStudy (anonymous):

Solve. 2x – 7 = 6x – 15 2 3 5 6

OpenStudy (anonymous):

x=2

OpenStudy (emmyma):

Simplifying 2x + -7 = 6x + -15 Reorder the terms: -7 + 2x = 6x + -15 Reorder the terms: -7 + 2x = -15 + 6x Solving -7 + 2x = -15 + 6x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-6x' to each side of the equation. -7 + 2x + -6x = -15 + 6x + -6x Combine like terms: 2x + -6x = -4x -7 + -4x = -15 + 6x + -6x Combine like terms: 6x + -6x = 0 -7 + -4x = -15 + 0 -7 + -4x = -15 Add '7' to each side of the equation. -7 + 7 + -4x = -15 + 7 Combine like terms: -7 + 7 = 0 0 + -4x = -15 + 7 -4x = -15 + 7 Combine like terms: -15 + 7 = -8 -4x = -8 Divide each side by '-4'. x = 2 Simplifying x = 2

OpenStudy (anonymous):

@parakeet is right x=2

OpenStudy (emmyma):

Thank you @HarukaN for the medal ^^~~

OpenStudy (anonymous):

poor parakeet never got the fluttering medal

OpenStudy (anonymous):

No problem! thank you for explaining so well! @emmyma

OpenStudy (anonymous):

i know :(

OpenStudy (anonymous):

I fanned you both :)

OpenStudy (anonymous):

yay!